OpenStudy (anonymous):

What is the prime factorization of 621?

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (anonymous):

start dividing by 2, then 3, then 5, then 7. etc until you cant divide no more.

OpenStudy (anonymous):

Since the sum of the given digits = 9 i.e. ( 6+2+1=9) so this number is divisible by 3 start dividing it by 3

OpenStudy (anonymous):

621 = 3 *207 = 3 *3* 69 = 3 *3* 3* 23 since 23 is the prime number, hence it can not be further factorised. Therefore factor od 621 is 3 *3* 3* 23 @maclipfo

OpenStudy (anonymous):

Thanks!